OpenStudy (anonymous):

Using interval notation, identify the domain for the function: f(x)= sqrt of x-9

OpenStudy (freckles):

\[f(x)=\sqrt{x}-9 ?\] just remember you don't want negatives under the square root part

OpenStudy (freckles):

so you want the thing inside the square root to be positive or zero if something is positive it is greater then 0 if something is zero it is equal to 0 so if something is positive or zero then it is >=0. so if the function really is f(x)=sqrt(x)-9 then the domain is easy to find if you meant f(x)=sqrt(x-9) then the domain is still easy to find you just have to do a little solving ...since remember you want x-9 to be positive or zero